OpenStudy (tal_tal):

The width of a rectangle is shown below: A coordinate plane with a point A at negative 3, 3 and C at negative 3, negative 4. If the area of the rectangle to be drawn is 35 square units, where should points B and D be located if they lie to the right of points A and C? B(2, 4) and D(2, −3) B(2, 3) and D(2, −4) B(3, 3) and D(3, −4) B(3, −4) and D(3, 3)

OpenStudy (tal_tal):

@mathmath333

OpenStudy (tal_tal):

@mathmath333 can you help please :)

OpenStudy (mathmath333):

length of AC=7 Area= l*b 35=7*b, b=5

OpenStudy (tal_tal):

so B?

OpenStudy (mathmath333):

https://www.desmos.com/calculator/bflhraq5nl

OpenStudy (mathmath333):

yep